OpenStudy (anonymous):

Choose the product. 4p^3(p^2 + 8p - 3)

OpenStudy (anonymous):

4p^5 + 12p^4 + 7p^3 4p^6 + 44p^3 4p^3 - 12p^2 +32p^4 4p^5 + 32p^4 -12p^3

OpenStudy (goformit100):

Correct

OpenStudy (anonymous):

what one

OpenStudy (goformit100):

it's c

OpenStudy (anonymous):

I thought it was D

OpenStudy (anonymous):

its definetly D

OpenStudy (callisto):

4p^3(p^2 + 8p - 3) = 4p^3(p^2) + 4p^3(8p) - 4p^3(3) = 4p^(3+2) + 8(4) p^(3+1) - 4(3) p^3 = ...? Can you try to do it?

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

when you are multiplying variables which have powers, you add the powers

OpenStudy (anonymous):

this might help you look it at it better, re writing P as X \[4x^3 ( x^2 + 8x + 3) \]

OpenStudy (anonymous):

sorry it should be a - 3 not a +3 up there

OpenStudy (anonymous):

oh ok

OpenStudy (callisto):

= 4p^(3+2) + 8(4) p^(3+1) - 4(3) p^3 what is 3+2 ?

OpenStudy (anonymous):

5

OpenStudy (callisto):

So, we get = 4p^5 + 8(4) p^(3+1) - 4(3) p^3 What is 8x4 ?

OpenStudy (anonymous):

32x

OpenStudy (goformit100):

8 x 4=32

OpenStudy (anonymous):

so its already got to be d

OpenStudy (callisto):

lol Yes :)

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome!